Compute the radius from the instantaneous center of rotation to the center of the vehicle rear axle (in m) required for an autonomous vehicle to follow the desired path based on the information below.

The lookahead distance is 10 m; the car length is 4 m; the angle between the vehicle’s body heading and the lookahead line is 30°. Your answer should be an integer.
1
Expert's answer
2020-12-07T03:46:48-0500

lookhead distance=10 m,length of car =4 m ,the angle between the vehicle’s body heading and the lookahead line is 30°



For I-Centre of rotation, here we know that R is the radius of from I-Centre which can be given by formula

"R=\frac{l_d}{2sin \alpha}=\frac{10}{2\times sin 30}= 10 m"

radius of Centre is 10 m